### v1

# Classification of Hoyle State Decay Branches in Active Target Time Projection Chamber using Neural Network

Abstract
A multi-class convolutional neural network (CNN) model has been developed using Keras deep learning library in Python for image-based classification of 12C Hoyle state decay branches from tracking information, recorded by Saha Active Target Time Projection Chamber, SAT-TPC (currently under development). The nuclear events, produced by the 30 MeV $\alpha$-particle beam in the SAT-TPC, filled with Ar + CO2 (90:10) gas mixture at atmospheric pressure, have been considered for training and validation of the models. The elastic scattering and Hoyle state sequential and direct decay events in the interaction of $\alpha$-particle with 40Ar, 12C, 16O nuclei have been generated through Monte-Carlo simulation. The three-dimensional tracks, produced by the scattering and decay products through primary ionization of gaseous medium, have been simulated with Geant4. The primary tracks, distributed on the beam-plane, have been convoluted with electron diffusion, obtained with Magboltz, to produce the final tracking information. The classification performance of the proposed model for different readout segmentation schemes of the SAT-TPC has been discussed.
1 Introduction
The formation of 12C by nuclear fusion of three $\alpha$-nuclei is a significant step in the stellar nucleo-synthesis. It plays an important role to produce the elements heavier than 4He and bridge the well-known gap, caused by the instability of nuclei of atomic weight 5 and 8. To explain the observed abundance of 12C in stellar media, the reaction ${}^{8}Be(\alpha,\gamma)^{12}C$ at the second step of triple-$\alpha$ fusion reaction was proposed to proceed through s-wave resonance around 7.68 MeV above the ground level in 12C [1]. An excited state with spin-parity $0^{+}$ and energy 7.65 MeV, namely the Hoyle state, was later verified experimentally [2]. Aside from its key role in the synthesis of the elements, there are several intriguing aspects about the structure of the Hoyle state. It has been theoretically conjectured as a molecule-like $\alpha$-cluster state [3, 4], or Bose-Einstein Condensate (BEC)-like configuration [5], etc. Eventually, many experimental explorations as well as theoretical modelling have followed suit to envisage the exotic structure of the Hoyle state [6]. In this context, investigation of the Hoyle state decay has turned out a relevant observable as the relative branching ratio of the direct decay of 12C to triple-$\alpha$ and the two-step sequential decay via the ground state of 8Be can be predicted following a formulation of the BEC wave function considering two and three-body tunnelling processes [7]. This investigation has an impact in understanding the stellar nucleosynthesis process as well. Any deviation from the sequential decay should modify the calculation of the nucleosynthesis reaction rate where exclusively the two-step sequential fusion is taken into account. Several experiments have been carried out to determine the branching ratio of the direct decay which has been found to be as small as 0.019% by one of the latest measurements [8]. The said limit has further been modified to 0.00057% (5.7 $\times$ 10-6) following indirect measurements [9]. These reports imply the necessity of measuring the cross-section of the Hoyle state decay branches to determine their branching ratio more precisely.
The Time Projection Chamber (TPC) [10, 11], equipped with long drift volume and position-sensitive electron multiplier readout, is a stand-alone device that can provide three-dimensional position information of charged particles. It utilizes the drift time together with the two-dimensional position information of the electrons, produced in the primary ionization of the gaseous molecules of the active medium by the passage of the particle to produce the three-dimensional position. In addition to wide usage in high energy physics experiments, eventually it has found application in low-energy nuclear physics experiments as well for measuring nuclear reaction cross-section. The relevant reaction kinematics can be studied by three-dimensional tracking of the reaction candidates (projectiles, ejectiles) using the same working principle of the TPC. However, the only change for these applications is implemented by opting the active gas medium as a target of the nuclear reaction. Thus, the device, known as Active Target Time Projection Chamber (AT-TPC) [12], can offer advantages to circumvent the issues related to the use of solid targets and elaborate detector arrangements. It may be further beneficial in evading the limitations of background, associated with the use of silicon detector arrays in measuring very low branching ratio, as pertinent to the exploration of the Hoyle state configuration. A high-sensitivity measurement of the direct decay mode has been performed recently using an AT-TPC [13]. By applying a Bayesian approach to study the contribution of the direct decay using a likelihood function, a limit of the direct decay mode less than 0.043% at the 95% confidence level has been achieved.
In this context, Machine Learning (ML)-based methods, which have shown significant improvements in classification problems [14], may be found useful in designating the relevant tracks and events. In particle physics experiments with an active target, such as MicroBooNE in the Short Baseline Neutrino program at Fermilab, a liquid argon TPC has incorporated ML methods to classify events and particle tracks [15]. It has also found application in low-energy nuclear physics experiments in recent times [16, 17]. The neural network technique offers a large class of computational models in ML methods, inspired by the way biological neural network in the human brain processes information. It consists of interconnected layers of nodes, also known as perceptrons. Each of these nodes processes input data, collected from others through non-linear communication, defined by weights. Subsequently, it passes the result, added with approximate biases, to the next layer. The network is trained by adjusting the weights and biases to minimize the error between its output and the recommended result for a given input. Several regularization techniques are used in training the network to prevent overfitting and improve generalization. While many different network configurations exist, certain models are found useful in classification problems. Convolutional Neural Network (CNN) is one of the specialised models that is designed to visualize grid-like structured data, such as one-dimensional time series or two-dimensional grid of pixels of an image [18]. The network employs a specialized kind of linear operation, called convolution, in place of general matrix multiplication in at least one of its layers. It is widely used in the field of computer vision for object detection and image recognition and known to involve fewer computational resources.
We have designed a prototype AT-TPC at Saha Institute of Nuclear Physics, namely the SAT-TPC [19], for studying Hoyle state decay mechanism and measure the branching ratio of the direct and sequential decay branches. The optimization of design and operational parameters of the SAT-TPC has been carried out with the help of numerical simulation and discussed in [19]. The present work reports the scheme of data analysis to be adopted for classification of Hoyle state decay events using a CNN model when the SAT-TPC, filled with an active gas mixture Ar + CO2 of volumetric ratio 90:10 at atmospheric pressure, would be subject to 30 MeV $\alpha$-particle beam. The relevant nuclear events, like elastic scattering and the Hoyle state decay following the inelastic scattering of the $\alpha$-projectile from the active target nuclei 40Ar, 12C and 16O have been generated following numerical simulation of non-relativistic nuclear interaction. The tracks, produced by the scattering and decay products in the horizontal plane containing the beam axis, referred to as the beam plane in this text, have been reconstructed following the primary ionization of active gaseous medium and transport of the primary electrons under the action of the electric field in the drift volume of the SAT-TPC. The set of correlated tracks for each event has been considered as a representative 2D image. A large number of such images, representing each class of events, has been produced in order to build the input dataset. A part of the dataset has been used for training and validation of the CNN model while the rest has been utilized for testing the model for its performance in classification of direct and sequential decay branches and their segregation from the elastic scattering events, termed hereafter as background. An optimization study of the readout granularity of the SAT-TPC, which in principle governs the image resolution, has been performed on the basis of classification efficacy of the CNN model. This approach has the potential to serve as an automated analysis framework for tagging and separating actual experimental data.
The article has been organized in the following manner. The workflow of the present numerical study starting from the generation of the scattering and decay events to the training of the CNN model has been described in section 2. All the stages of the numerical simulation have been briefly discussed in several sub-sections of the same. The results of the implementation of the CNN model in classification of nuclear events have been presented in section 3. The article concludes with section 4.
2 Numerical Study
The flowchart of the numerical model followed in this work to address the classification problem of the Hoyle state decay branches in the case of 30 MeV $\alpha$-projectile interacting with Ar + CO2 (90:10) gas mixture, is illustrated in figure 1. The relevant software toolkits associated with each stage to perform the numerical work are mentioned as well. A brief discussion on individual stages has been provided in the following sub-sections.
Figure 1: Flowchart of the numerical model
2.1 Event Generation
In the interaction of 30 MeV $\alpha$-projectile with the active gas mixture Ar + CO2, five cases of nuclear events have been considered in this work. These are the elastic scattering of the $\alpha$-projectile with the active targets 40Ar, 12C, 16O nuclei and the inelastic scattering of the $\alpha$-particle with 12C only, followed by direct and sequential decay of the Hoyle state, each producing three $\alpha$-particles.
An event-by-event Monte Carlo simulation of the elastic scattering has been developed in ROOT framework [20] using the TGenPhaseSpace class following non-relativistic kinematics for all possible angles in the center of mass (c.m.) frame. The energy and corresponding angle of the scattering products have been obtained by boost in the laboratory (lab) frame. The correlation of the angle, $\theta$, and the energy in the lab frame of the products of the three cases of elastic scattering $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O, are plotted in the figure 2, respectively. These data have been used to produce the primary ionization tracks of the corresponding scattering products, as discussed in sub-section 2.2.
(a) $\alpha$ + 40Ar
(b) $\alpha$ + 12C
(c) $\alpha$ + 16O
Figure 2: Correlated angle and energy of scattering products in lab frame for elastic scattering cases: (a) $\alpha$ + 40Ar, (b) $\alpha$ + 12C, and (c) $\alpha$ + 16O
We have investigated the dynamics of the Hoyle state focusing on the excitation of 12C nucleus through inelastic scattering of $\alpha$-particle followed by the decay mechanisms. The direct decay of the 12C as well as its sequential decay via 8Be + $\alpha$, all leading to triple-$\alpha$ products, have been modeled using a phase-space generator. The resulting energy, momentum and angular distribution of the $\alpha$-particles have been obtained after applying a relativistic boost to the decay products of the Hoyle state, generated in the c.m. frame. For the direct decay of the Hoyle state, the triple-$\alpha$ decay has been boosted with the four-momentum vector of 12C. In case of sequential decay, it has been done in two steps. In the first step, $\alpha$ and 8Be have been generated and boosted in the direction of 12C. In the next step, the two $\alpha$-particles, generated from 8Be, have been boosted in the direction of 8Be. The four-momentum, obtained for all the nuclei in each event, has been used to produce the corresponding tracks, as detailed in the next sub-section 2.2. These parameters have been recorded in histograms and the 2D correlation plots of the angle, $\theta$, and the energy in the lab frame, are shown in figure 3.
(a) Direct decay
(b) Sequential decay
Figure 3: Correlated angle and energy in the lab frame of three $\alpha$-particles in Hoyle state decay branches: (a) Direct decay, and (b) Sequential decay
2.2 Track Reconstruction
The geometry of the SAT-TPC, following the schematic design which is depicted in figure 4(a), has been modeled in Geant4 [21]. The uniform drift field of 500 V/cm in the active region of the TPC along the z axis has been considered for this work.
The primary ionization in the active gas volume, caused by the scattering and decay products which are ejected with respective energy along their trajectories within the active gaseous volume of the SAT-TPC, has been simulated with Geant4. To incorporate the low-energy electromagnetic processes, the physics lists Penelope, Livermore, and Photo Absorption and Ionization (PAI) in Geant4, have been taken into account. Only those events with their product tracks, lying on the beam plane axis in the active volume of the SAT-TPC, have been taken into account to simplify the calculation. For each case of elastic scattering and Hoyle state decay branches,1000 such events have been considered (i.e. 5000 in total). The correlated primary tracks of all the cases of the nuclear events, as obtained from Geant4, have been collectively illustrated in figure 4(b) for 100 events.
Figure 4: (a) Schematic design of the SAT-TPC, (b) Correlated primary tracks of 100 events of each class, as simulated in Geant4
In the next step, the primary tracks for each event have been projected on the readout plane of the SAT-TPC with sensitive area 12 cm $\times$ 12 cm and consisting of 60 $\times$ 60 square pixels. No temporal evolution of the primary tracks through the drift volume of the SAT-TPC has been considered as the simulation of the same for 5000 events would be computationally expensive. However, the effect of diffusion has been convoluted in the primary tracks using a Gaussian profile with a variance equivalent to the transverse diffusion of the electrons in the given electric field of the SAT-TPC, as produced by the Magboltz toolkit [22] in the Garfield++ framework [23]. A detailed discussion on the temporal evolution of the primary tracks may be found in our previous work [19]. Three typical examples of reconstructed tracks, projected on the readout plane with 60 $\times$ 60 pixels, each for three cases of elastic scattering, are illustrated in figure 5. It can be observed that the reconstructed track of the scattered 40Ar is too short in comparison to that of the scattered 12C and 16O nuclei, being the heaviest among the three products. Similarly, three examples of the reconstructed tracks for each of the direct and sequential decay events, following convolution and projection of the primary tracks on the readout plane, are shown in figure 6. These reconstructed tracks corresponding to each class of nuclear events have been recorded as input image data for training and validation, followed by testing of the performance of the CNN model in classification of Hoyle state decay branches.
(a) $\alpha$ + 40Ar
(b) $\alpha$ + 12C
(c) $\alpha$ + 16O
Figure 5: Reconstructed tracks on the SAT-TPC readout plane of the elastic scattering cases: (a) $\alpha$ + 40Ar, (b) $\alpha$ + 12C, and (c) $\alpha$ + 16O
(a) Direct decay
(b) Sequential decay
Figure 6: Reconstructed tracks on the SAT-TPC readout plane of the Hoyle state decay branches: (a) Direct decay, and (b) Sequential decay
2.3 CNN Model
In the first step, a convolution filter having smaller width and height and the same depth as that of input volume is used to extract relevant features from the input data, such as edges, corners, shapes, etc. The filter slides over the input grid performing element-wise multiplication and summing to produce two-dimensional results without changing the dimension. The results produced by each filter are stacked together, and as a result, we get output volume having a depth equal to the number of filters. In the next step, activation layer adds nonlinearity to the network by applying an element-wise activation function to the output of the convolution layer, keeping the data volume unchanged. It allows the network to learn more complex patterns and a few examples of the same are ReLU (Rectified Linear Unit), Tanh, etc. After this layer, a pooling function is used to down-sample the input data of the previous layer, which in turn helps in reducing computational load, and thus improving the efficiency of the network. It also helps to minimize overfitting, Pooling reduces dimensionality, and helps prevent overfitting. Additionally, dropout is used to reduce the over-reliance on specific neurons. Common types of pooling function include max pooling and average pooling. It is followed by the flattening of the resulted feature map into a one-dimensional vector. In the fully connected layer, the final classification or regression task is performed on the basis of the learned features. At the end stage, the output from the fully connected layers is fed into a logistic function for classification to convert the output into the probability score of each class, here the elastic scattering and the Hoyle state decay events.
In the present work, we have used the VGG-16 (Visual Geometry Group) architecture [24] to design the CNN model which is trained on ImageNet repository [25]. It is a widely popular and highly efficient CNN architecture for image recognition and very competitive even today, compared to other architectures, such as ResNet or DenseNet. The VGG-16 is well known for its simple structure and distinguished by the application of small convolution filter of size 3 $\times$ 3 pixels. The scheme of the said architecture, as has been used in the current work for classification of the nuclear events, is illustrated in figure 7. The input shape for the CNN model is [channels, height, width] which is equivalent to the two-dimensional color images with red, green and blue values in the third dimension of the image pixels. For our study, we have reshaped the input layer of VGG-16 architecture into [60,60,3] shape and the model receives input as 60 $\times$ 60 pixel event images with three channels to leverage pre-trained ImageNet weights. There are five pooling layers in the model as per the architecture or five blocks with one pooling layer. The first two convolutional blocks have 2 convolutional layers, and the later ones have 3 convolutional layers each. Early layers deal with large images, so fewer layers are used to reduce computation. As the image size shrinks, more layers are added to capture complex features. This design builds depth gradually while keeping the model efficient and powerful. For the fully connected CNN model, we have utilized a single hidden layer architecture, created with Keras, a high-level deep learning library in Python [26]. Keras acts as a user-friendly interface for TensorFlow which is a library that facilitates efficient numerical computation and automatic differentiation. The output layer consists of five nodes, representing event classification with integer labels 0, 1, 2, 3 and 4, representing the direct, sequential decay and $\alpha$ + 40Ar, $\alpha$ + 12C and $\alpha$ + 16O scattering events, respectively, as shown in figure 7.
Figure 7: Scheme of VGG-16 architecture in the present CNN model
[…]
Epochs 20
Batch size 128
Learning rate 0.001
Loss categorical crossentropy
Optimizer Adam
Table 1: CNN model parameters and values
[…]
3 Performance Analysis
In this section, the evaluation of the training and validation, followed by classification performance of the model, have been discussed. The effect of variation in readout segmentation of the SAT-TPC has also been studied.
3.1 Training & Validation Accuracy
To evaluate the training of the present CNN model using regularization technique, the accuracy and loss curves of training and validation have been produced, as shown in figure 8. The curves show that the performance of the model has improved as the number of epoch has increased and stabilized after 10.
Figure 8: (a) Accuracy, and (b) Loss of the training and validation processes
3.2 Classification Performance
We have evaluated the performance of the CNN model using three key metrics: precision, recall, and F1 score. These metrics have provided more detailed insights into its performance and errors than accuracy alone. For the track classification, we have focused on the ability of the model to identify direct and sequential decay events and the backgrounds accurately. The precision is a measure of the accuracy of a model in identifying the positive instances which is calculated as the ratio of correctly predicted positive observations to the total predicted positives. In simpler terms, it tells us how many of the items that the model identifies as positive are actually positive. Recall, also known as sensitivity, measures the ability of the model to identify all relevant instances within a dataset. It is calculated as the ratio of correctly predicted positive observations to all observations in the actual class. Essentially, it indicates how many of the actual positives the model is able to capture. The F1 score is the harmonic mean of precision and recall. It provides a single metric that balances both the precision and recall of a model, especially useful when the classes are imbalanced. The F1 score combines the strengths of precision and recall, making it a comprehensive metric to evaluate the performance of a model. It is useful particularly in situations where the cost of false positives and false negatives is significant. The support column indicates the number of true instances (actual samples) of each event class in the dataset, used to evaluate the model. The performance metrics of the present model are given in table 2. These have been determined with a readout segmentation of 2 mm $\times$ 2 mm.
Table 2: Classification report of VGG-16 based CNN model
Readout segmentation Event classification task Precision Recall F1 Support
2 mm $\times$ 2 mm direct decay 0.70 0.65 0.67 200
sequential decay 0.68 0.72 0.70 200
$\alpha$+40Ar scattering 0.98 0.98 0.98 200
$\alpha$+12C scattering 0.94 0.94 0.94 200
$\alpha$+16O scattering 0.92 0.94 0.93 200
The two metrics, precision and recall, are best understood by producing a confusion matrix as shown in figure 9 which facilitates assessment of the performance and errors of the CNN model. It is evident from the matrix that the model has exhibited strong performance in classifying the scattering events $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O, with high accuracy. The matrix reveals the misclassifications between the direct and sequential decays which indicates that these classes might require more distinct training data. In order to test this concept, the readout segmentation has been varied to study its effect on the performance of the model (discussed in the next sub-section).
Figure 9: Confusion matrix
3.3 Effect of Readout Segmentation
The scheme of the anode plane segmentation of the SAT-TPC in the simulation model has been varied to 1 mm $\times$ 1 mm and 3 mm $\times$ 3 mm pixels. The calculation has been carried out with the same class of events, discussed in the previous section. The classification performance for the two segmentation schemes in terms of the three metrics is given in table 3. It shows that the reduction in the size of the readout pixel from 3 mm $\times$ 3 mm to 1 mm $\times$ 1 mm has led to improved classification accuracy, particularly for decay events. This suggests that finer segmentation enhances spatial resolution, allowing the CNN to distinguish subtle differences in track morphology.
Table 3: Classification report for segmentation schemes 1 mm $\times$ 1 mm and 3 mm $\times$ 3 mm
readout segmentation event classification task Precision Recall F1 Support
1 mm $\times$ 1 mm direct decay 0.62 0.67 0.64 200
sequential decay 0.64 0.57 0.61 200
$\alpha$+40Ar scattering 0.94 0.90 0.92 200
$\alpha$+12C scattering 0.86 0.85 0.86 200
$\alpha$+16O scattering 0.80 0.86 0.83 200
3 mm $\times$ 3 mm direct decay 0.90 0.82 0.86 200
sequential decay 0.83 0.90 0.87 200
$\alpha$+40Ar scattering 0.99 1.00 1.00 200
$\alpha$+12C 0.99 1.00 0.99 200
$\alpha$+16O scattering 0.99 0.98 0.99 200
The corresponding confusion matrices, as determined for two different readout segmentation schemes, are shown in figure 10. The results clearly indicate that the readout segmentation 1 mm $\times$ 1 mm provides better resolution with respect to other choices. However, the improvement in the accuracy of event classification has been reflected mostly in the case of distinguishing direct and sequential decay events.
(a) 3 mm x 3 mm pixels
(b) 1 mm x 1 mm pixels
Figure 10: Comparison of confusion matrices for different pixel sizes
4 Summary & Conclusion
The classification of Hoyle state decay events and the background elastic scattering events forms the central theme of the present study, showcasing the application of supervised machine learning, particularly CNN, for accurate segregation of the relevant class of events. The work has highlighted the challenges of distinguishing between direct decay, sequential decay, and the background scattering events $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O in the context of using 30 MeV $\alpha$-beam and a prototype SAT-TPC, filled with active gas target Ar + CO2. We have performed Monte Carlo simulations to study the nuclear interactions leading to formation of the Hoyle state, following excitation through inelastic scattering of $\alpha$ from the active target nucleus 12C and the relevant elastic scattering of $\alpha$ from other target nuclei 40Ar, 12C and 16O. The decay kinematics have been analyzed in a boosted frame, allowing the extraction of energy and angular distributions of the emitted $\alpha$-particles. The same has been done for the elastic scattering events to study the energy and angular distributions of the scattered products. The primary ionization by the beam and the scattering and decay products has been simulated in Geant4 using a low-energy physics list that has been used to reconstruct the event tracks in the SAT-TPC active volume, convoluted with diffusion parameter and projected on two-dimensional readout with segmentation 2 mm $\times$ 2 mm. Using the VGG-16 CNN architecture, the model has classified the events on the basis of the track morphology images. Regularization technique, such as dropout, has been employed to prevent overfitting and enhance the robustness of the model, achieving high precision, recall, and F1 scores across all categories.
Despite strong classification performance, the study has identified specific trends in misclassifications, particularly among the decay events. The classification performance of the model for the decay events has been found to improve with the finer readout segmentation as it has reduced the misclassification rate of the decay events. These results have proved the efficiency of the model in distinguishing Hoyle state decays from the background events while highlighting areas for further optimization, such as refining feature representations and detector configurations to enhance separation between closely related scattering categories.
This work has demonstrated the utilization of CNN in analyzing experimental data of low-energy nuclear physics experiments performed with SAT-TPC, providing a framework for advancing robust event classification in nuclear and astrophysical research. Future work will explore advanced architectures, such as ResNet, and transformer-based models, to further enhance classification performance. In addition, real-time processing capabilities will be investigated to enable integration with experimental data acquisition systems.
Acknowledgments
The authors acknowledge the support provided by their institute. They are thankful to Dr. Tapan Kumar Rana and Dr. Santu Manna of the Variable Energy Cyclotron Center, and Prof. Satyaki Bhattacharya of Saha Institute of Nuclear Physics Kolkata, India, for valuable discussion and suggestions.

### v2

@@
 Abstract
-A multi-class convolutional neural network (CNN) model has been developed using Keras deep learning library in Python for image-based classification of 12C Hoyle state decay branches from tracking information, recorded by Saha Active Target Time Projection Chamber, SAT-TPC (currently under development).
-The nuclear events, produced by the 30 MeV $\alpha$-particle beam in the SAT-TPC, filled with Ar + CO2 (90:10) gas mixture at atmospheric pressure, have been considered for training and validation of the models.
-The elastic scattering and Hoyle state sequential and direct decay events in the interaction of $\alpha$-particle with 40Ar, 12C, 16O nuclei have been generated through Monte-Carlo simulation.
-The three-dimensional tracks, produced by the scattering and decay products through primary ionization of gaseous medium, have been simulated with Geant4.
+A multi-class convolutional neural network model has been developed using the Keras deep learning library in Python for image-based classification of 12C Hoyle state decay branches from tracking information, to be recorded by a prototype Active Target Time Projection Chamber, SAT-TPC (currently under development), filled with an Ar + CO2 (90:10) gas mixture at atmospheric pressure.
+The elastic scattering and Hoyle state sequential and direct decay events, in the interaction of 30 MeV $\alpha$-particle beam with 40Ar, 12C, 16O nuclei, have been generated through Monte-Carlo simulation.
+The three-dimensional tracks, produced by the scattering and decay products through primary ionization of the gaseous medium, have been simulated with Geant4.
 The primary tracks, distributed on the beam-plane, have been convoluted with electron diffusion, obtained with Magboltz, to produce the final tracking information.
@@
 The formation of 12C by nuclear fusion of three $\alpha$-nuclei is a significant step in the stellar nucleo-synthesis.
-It plays an important role to produce the elements heavier than 4He and bridge the well-known gap, caused by the instability of nuclei of atomic weight 5 and 8.
-To explain the observed abundance of 12C in stellar media, the reaction ${}^{8}Be(\alpha,\gamma)^{12}C$ at the second step of triple-$\alpha$ fusion reaction was proposed to proceed through s-wave resonance around 7.68 MeV above the ground level in 12C [1].
+It plays an important role in producing the elements heavier than 4He and bridge the well-known gap, caused by the instability of nuclei of atomic weight 5 and 8.
+To explain the observed abundance of 12C in stellar media, the reaction $^{8}Be(\alpha,\gamma)^{12}C$ at the second step of the triple-$\alpha$ fusion reaction was proposed to proceed through s-wave resonance around 7.68 MeV above the ground level in 12C [1].
 An excited state with spin-parity $0^{+}$ and energy 7.65 MeV, namely the Hoyle state, was later verified experimentally [2].
@@
 Eventually, many experimental explorations as well as theoretical modelling have followed suit to envisage the exotic structure of the Hoyle state [6].
-In this context, investigation of the Hoyle state decay has turned out a relevant observable as the relative branching ratio of the direct decay of 12C to triple-$\alpha$ and the two-step sequential decay via the ground state of 8Be can be predicted following a formulation of the BEC wave function considering two and three-body tunnelling processes [7].
+In this context, the decay modes of the Hoyle state serve as valuable observables.
+The relative branching ratio between the direct triple-$\alpha$ decay of 12C and the sequential two-step decay via the ground state of 8Be can be theoretically estimated using a BEC-based wave function that includes two and three-body tunnelling processes [7].
 This investigation has an impact in understanding the stellar nucleosynthesis process as well.
@@
 These reports imply the necessity of measuring the cross-section of the Hoyle state decay branches to determine their branching ratio more precisely.
-The Time Projection Chamber (TPC) [10, 11], equipped with long drift volume and position-sensitive electron multiplier readout, is a stand-alone device that can provide three-dimensional position information of charged particles.
-It utilizes the drift time together with the two-dimensional position information of the electrons, produced in the primary ionization of the gaseous molecules of the active medium by the passage of the particle to produce the three-dimensional position.
-In addition to wide usage in high energy physics experiments, eventually it has found application in low-energy nuclear physics experiments as well for measuring nuclear reaction cross-section.
-The relevant reaction kinematics can be studied by three-dimensional tracking of the reaction candidates (projectiles, ejectiles) using the same working principle of the TPC.
-However, the only change for these applications is implemented by opting the active gas medium as a target of the nuclear reaction.
-Thus, the device, known as Active Target Time Projection Chamber (AT-TPC) [12], can offer advantages to circumvent the issues related to the use of solid targets and elaborate detector arrangements.
-It may be further beneficial in evading the limitations of background, associated with the use of silicon detector arrays in measuring very low branching ratio, as pertinent to the exploration of the Hoyle state configuration.
+The Time Projection Chamber (TPC) [10, 11], equipped with a long drift volume and position-sensitive electron multiplier readout, is a stand-alone device that can provide three-dimensional (3D) position information of charged particles.
+It utilizes the drift time together with the two-dimensional (2D) position information of the electrons, produced in the primary ionization of the gaseous molecules of the active medium by the passage of the particle to produce the 3D position.
+In addition to wide usage in high-energy particle physics experiments, eventually it has found application in low-energy nuclear physics experiments as well for measuring nuclear scattering cross-sections.
+The relevant kinematics can be studied by 3D tracking of the scattering products using the same working principle of the TPC.
+However, the only change for these applications is implemented by opting the active gas medium as the target nucleus.
+Thus, the device, known as an Active Target Time Projection Chamber (AT-TPC) [12], can offer advantages to circumvent the issues related to the use of solid targets and elaborate detector arrangements.
+It may be further beneficial in evading the limitations of background, associated with the use of silicon detector arrays in measuring very low branching ratios, as pertinent to the exploration of the Hoyle state configuration.
 A high-sensitivity measurement of the direct decay mode has been performed recently using an AT-TPC [13].
@@
 In particle physics experiments with an active target, such as MicroBooNE in the Short Baseline Neutrino program at Fermilab, a liquid argon TPC has incorporated ML methods to classify events and particle tracks [15].
-It has also found application in low-energy nuclear physics experiments in recent times [16, 17].
-The neural network technique offers a large class of computational models in ML methods, inspired by the way biological neural network in the human brain processes information.
+It has also found applications in nuclear physics experiments in recent times [16, 17].
+The neural network technique offers a large class of computational models in ML methods, inspired by the way the biological neural network in the human brain processes information.
 It consists of interconnected layers of nodes, also known as perceptrons.
 Each of these nodes processes input data, collected from others through non-linear communication, defined by weights.
-Subsequently, it passes the result, added with approximate biases, to the next layer.
+Subsequently, it passes the results, added with approximate biases, to the next layer.
 The network is trained by adjusting the weights and biases to minimize the error between its output and the recommended result for a given input.
@@
 While many different network configurations exist, certain models are found useful in classification problems.
-Convolutional Neural Network (CNN) is one of the specialised models that is designed to visualize grid-like structured data, such as one-dimensional time series or two-dimensional grid of pixels of an image [18].
+Convolutional Neural Network (CNN) is one of the specialised models that is designed to visualize grid-like structured data, such as one-dimensional (1D) time series or 2D grids of pixels of an image [18].
 The network employs a specialized kind of linear operation, called convolution, in place of general matrix multiplication in at least one of its layers.
 It is widely used in the field of computer vision for object detection and image recognition and known to involve fewer computational resources.
-We have designed a prototype AT-TPC at Saha Institute of Nuclear Physics, namely the SAT-TPC [19], for studying Hoyle state decay mechanism and measure the branching ratio of the direct and sequential decay branches.
+We have designed a prototype AT-TPC at Saha Institute of Nuclear Physics, namely the SAT-TPC [19], to study the Hoyle state decay mechanism and measure the branching ratio of the direct and sequential decay branches.
 The optimization of design and operational parameters of the SAT-TPC has been carried out with the help of numerical simulation and discussed in [19].
 The present work reports the scheme of data analysis to be adopted for classification of Hoyle state decay events using a CNN model when the SAT-TPC, filled with an active gas mixture Ar + CO2 of volumetric ratio 90:10 at atmospheric pressure, would be subject to 30 MeV $\alpha$-particle beam.
-The relevant nuclear events, like elastic scattering and the Hoyle state decay following the inelastic scattering of the $\alpha$-projectile from the active target nuclei 40Ar, 12C and 16O have been generated following numerical simulation of non-relativistic nuclear interaction.
-The tracks, produced by the scattering and decay products in the horizontal plane containing the beam axis, referred to as the beam plane in this text, have been reconstructed following the primary ionization of active gaseous medium and transport of the primary electrons under the action of the electric field in the drift volume of the SAT-TPC.
+The relevant nuclear events, like elastic scattering and the Hoyle state decay following the inelastic scattering of the $\alpha$-projectile from the active target nuclei 40Ar, 12C and 16O, have been generated following numerical simulation of non-relativistic nuclear interaction.
+The tracks, produced by the scattering and decay products in the horizontal plane containing the beam axis, (referred to as the beam plane in this text), have been reconstructed following the primary ionization of the active gaseous medium and transport of the primary electrons under the action of the electric field in the drift volume of the SAT-TPC.
 The set of correlated tracks for each event has been considered as a representative 2D image.
@@
 2 Numerical Study
-The flowchart of the numerical model followed in this work to address the classification problem of the Hoyle state decay branches in the case of 30 MeV $\alpha$-projectile interacting with Ar + CO2 (90:10) gas mixture, is illustrated in figure 1.
+The flowchart of the numerical model, followed in this work to address the classification problem of the Hoyle state decay branches in the case of a 30 MeV $\alpha$-projectile interacting with an Ar + CO2 (90:10) gas mixture, is illustrated in figure 1.
 The relevant software toolkits associated with each stage to perform the numerical work are mentioned as well.
@@
 The energy and corresponding angle of the scattering products have been obtained by boost in the laboratory (lab) frame.
-The correlation of the angle, $\theta$, and the energy in the lab frame of the products of the three cases of elastic scattering $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O, are plotted in the figure 2, respectively.
+The correlation of the angle, $\theta$, and the energy in the lab frame, $E_{lab}$ of the products of the three cases of elastic scattering $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O, are plotted in the figure 2, respectively.
 These data have been used to produce the primary ionization tracks of the corresponding scattering products, as discussed in sub-section 2.2.
@@
 Figure 2: Correlated angle and energy of scattering products in lab frame for elastic scattering cases: (a) $\alpha$ + 40Ar, (b) $\alpha$ + 12C, and (c) $\alpha$ + 16O
-We have investigated the dynamics of the Hoyle state focusing on the excitation of 12C nucleus through inelastic scattering of $\alpha$-particle followed by the decay mechanisms.
+We have investigated the dynamics of the Hoyle state focusing on the excitation of the 12C nucleus through inelastic scattering of an $\alpha$-particle followed by the decay mechanisms.
 The direct decay of the 12C as well as its sequential decay via 8Be + $\alpha$, all leading to triple-$\alpha$ products, have been modeled using a phase-space generator.
@@
 The four-momentum, obtained for all the nuclei in each event, has been used to produce the corresponding tracks, as detailed in the next sub-section 2.2.
-These parameters have been recorded in histograms and the 2D correlation plots of the angle, $\theta$, and the energy in the lab frame, are shown in figure 3.
+These parameters have been recorded in histograms and the 2D correlation plots of the angle, $\theta$, and the energy $E_{lab}$ in the lab frame, are shown in figure 3.
 (a) Direct decay
 (b) Sequential decay
-Figure 3: Correlated angle and energy in the lab frame of three $\alpha$-particles in Hoyle state decay branches: (a) Direct decay, and (b) Sequential decay
+Figure 3: Correlated angle $\theta$ and energy $E_{lab}$ in the lab frame of three $\alpha$-particles in Hoyle state decay branches: (a) Direct decay, and (b) Sequential decay
 2.2 Track Reconstruction
-The geometry of the SAT-TPC, following the schematic design which is depicted in figure 4(a), has been modeled in Geant4 [21].
-The uniform drift field of 500 V/cm in the active region of the TPC along the z axis has been considered for this work.
-The primary ionization in the active gas volume, caused by the scattering and decay products which are ejected with respective energy along their trajectories within the active gaseous volume of the SAT-TPC, has been simulated with Geant4.
-To incorporate the low-energy electromagnetic processes, the physics lists Penelope, Livermore, and Photo Absorption and Ionization (PAI) in Geant4, have been taken into account.
-Only those events with their product tracks, lying on the beam plane axis in the active volume of the SAT-TPC, have been taken into account to simplify the calculation.
+The geometry of the SAT-TPC, depicted in figure 4(a), has been modeled in Geant4 [21].
+The uniform drift field of 500 V/cm in the active region of the TPC along the Z-axis has been considered for this work.
+The primary ionization in the active gas volume, caused by the scattering and decay products which are ejected with their respective energies along their trajectories within the active gaseous volume of the SAT-TPC, has been simulated with Geant4.
+To incorporate the low-energy electromagnetic processes, the physics lists Penelope, Livermore, and Photo Absorption and Ionization (PAI) in Geant4 have been taken into account.
+Only those events with their product tracks, lying on the beam plane (XY) in the active volume of the SAT-TPC, have been taken into account to simplify the calculation.
 For each case of elastic scattering and Hoyle state decay branches,1000 such events have been considered (i.e.
 5000 in total).
-The correlated primary tracks of all the cases of the nuclear events, as obtained from Geant4, have been collectively illustrated in figure 4(b) for 100 events.
-Figure 4: (a) Schematic design of the SAT-TPC, (b) Correlated primary tracks of 100 events of each class, as simulated in Geant4
-In the next step, the primary tracks for each event have been projected on the readout plane of the SAT-TPC with sensitive area 12 cm $\times$ 12 cm and consisting of 60 $\times$ 60 square pixels.
+The correlated primary tracks of all the cases of the nuclear events, as obtained from Geant4, have been collectively illustrated in figure 4(b) for 100 events in total with 20 for each case.
+Figure 4: (a) Schematic design of the SAT-TPC, and (b) correlated primary tracks of 100 events of each class, as simulated in Geant4
+In the next step, the primary tracks for each event have been projected on the readout plane of the SAT-TPC with sensitive area 12 cm $\times$ 12 cm, consisting of 60 $\times$ 60 square pixels.
 No temporal evolution of the primary tracks through the drift volume of the SAT-TPC has been considered as the simulation of the same for 5000 events would be computationally expensive.
-However, the effect of diffusion has been convoluted in the primary tracks using a Gaussian profile with a variance equivalent to the transverse diffusion of the electrons in the given electric field of the SAT-TPC, as produced by the Magboltz toolkit [22] in the Garfield++ framework [23].
+However, the effect of diffusion has been convoluted in the primary tracks using a Gaussian profile with a variance, equivalent to the transverse diffusion of the electrons in the given electric field of the SAT-TPC, as produced by the Magboltz toolkit [22] in the Garfield++ framework [23].
 A detailed discussion on the temporal evolution of the primary tracks may be found in our previous work [19].
@@
 2.3 CNN Model
-In the first step, a convolution filter having smaller width and height and the same depth as that of input volume is used to extract relevant features from the input data, such as edges, corners, shapes, etc.
-The filter slides over the input grid performing element-wise multiplication and summing to produce two-dimensional results without changing the dimension.
-The results produced by each filter are stacked together, and as a result, we get output volume having a depth equal to the number of filters.
-In the next step, activation layer adds nonlinearity to the network by applying an element-wise activation function to the output of the convolution layer, keeping the data volume unchanged.
+In the first step, a convolution filter having a smaller width and height and the same depth as the input volume, is used to extract relevant features from the input data, such as edges, corners, shapes, etc.
+The filter slides over the input grid performing element-wise multiplication and summing to produce 2D results without changing the dimension.
+The results, produced by each filter, are stacked together, and as a result, we get an output volume, having a depth equal to the number of filters.
+In the next step, an activation layer adds nonlinearity to the network by applying an element-wise activation function to the output of the convolution layer, keeping the data volume unchanged.
 It allows the network to learn more complex patterns and a few examples of the same are ReLU (Rectified Linear Unit), Tanh, etc.
 After this layer, a pooling function is used to down-sample the input data of the previous layer, which in turn helps in reducing computational load, and thus improving the efficiency of the network.
-It also helps to minimize overfitting, Pooling reduces dimensionality, and helps prevent overfitting.
+It also helps to minimize overfitting and reduce dimensionality.
 Additionally, dropout is used to reduce the over-reliance on specific neurons.
-Common types of pooling function include max pooling and average pooling.
-It is followed by the flattening of the resulted feature map into a one-dimensional vector.
+Common types of pooling functions include max pooling and average pooling.
+It is followed by the flattening of the resulting feature map into a 1D vector.
 In the fully connected layer, the final classification or regression task is performed on the basis of the learned features.
-At the end stage, the output from the fully connected layers is fed into a logistic function for classification to convert the output into the probability score of each class, here the elastic scattering and the Hoyle state decay events.
-In the present work, we have used the VGG-16 (Visual Geometry Group) architecture [24] to design the CNN model which is trained on ImageNet repository [25].
+At the end stage, the output from the fully connected layer is fed into a logistic function to convert the output into the probability score of each class, (the elastic scattering and the Hoyle state decay events).
+In the present work, we have used the VGG-16 (Visual Geometry Group) architecture [24] to design the CNN model which is trained on the ImageNet repository [25].
 It is a widely popular and highly efficient CNN architecture for image recognition and very competitive even today, compared to other architectures, such as ResNet or DenseNet.
-The VGG-16 is well known for its simple structure and distinguished by the application of small convolution filter of size 3 $\times$ 3 pixels.
+The VGG-16 is well known for its simple structure and distinguished by the application of a small convolution filter of size 3 $\times$ 3 pixels.
 The scheme of the said architecture, as has been used in the current work for classification of the nuclear events, is illustrated in figure 7.
-The input shape for the CNN model is [channels, height, width] which is equivalent to the two-dimensional color images with red, green and blue values in the third dimension of the image pixels.
-For our study, we have reshaped the input layer of VGG-16 architecture into [60,60,3] shape and the model receives input as 60 $\times$ 60 pixel event images with three channels to leverage pre-trained ImageNet weights.
+The input shape for the CNN model is [channels, height, width] which is equivalent to the 2D color images with red, green and blue values in the third dimension of the image pixels.
+For our study, we have reshaped the input layer of the VGG-16 architecture to a shape of [60,60,3] and the model receives input as 60 $\times$ 60 pixel event images with three channels to leverage pre-trained ImageNet weights.
 There are five pooling layers in the model as per the architecture or five blocks with one pooling layer.
-The first two convolutional blocks have 2 convolutional layers, and the later ones have 3 convolutional layers each.
+The first two convolutional blocks have two convolutional layers, and the later ones have three convolutional layers each.
 Early layers deal with large images, so fewer layers are used to reduce computation.
@@
 Keras acts as a user-friendly interface for TensorFlow which is a library that facilitates efficient numerical computation and automatic differentiation.
-The output layer consists of five nodes, representing event classification with integer labels 0, 1, 2, 3 and 4, representing the direct, sequential decay and $\alpha$ + 40Ar, $\alpha$ + 12C and $\alpha$ + 16O scattering events, respectively, as shown in figure 7.
+The output layer consists of five nodes, representing event classification with integer labels 0, 1, 2, 3 and 4, representing the direct, sequential decay branches and $\alpha$ + 40Ar, $\alpha$ + 12C and $\alpha$ + 16O scattering events, respectively, as shown in figure 7.
 Figure 7: Scheme of VGG-16 architecture in the present CNN model
@@
 3 Performance Analysis
-In this section, the evaluation of the training and validation, followed by classification performance of the model, have been discussed.
+In this section, the evaluation of the training and validation, followed by classification performance of the model, are discussed.
 The effect of variation in readout segmentation of the SAT-TPC has also been studied.
 3.1 Training & Validation Accuracy
-To evaluate the training of the present CNN model using regularization technique, the accuracy and loss curves of training and validation have been produced, as shown in figure 8.
-The curves show that the performance of the model has improved as the number of epoch has increased and stabilized after 10.
+To evaluate the training of the present CNN model using a regularization technique, the accuracy and loss curves of training and validation have been produced, as shown in figure 8.
+The curves show that the performance of the model has improved as the number of epochs has increased and stabilized after 10.
 Figure 8: (a) Accuracy, and (b) loss of the training and validation processes
@@
 It is useful particularly in situations where the cost of false positives and false negatives is significant.
-The support column indicates the number of true instances (actual samples) of each event class in the dataset, used to evaluate the model.
+The support column indicates the number of true instances (actual samples) of each event class in the dataset used to evaluate the model.
 The performance metrics of the present model are given in table 2.
@@
 $\alpha$+16O scattering 0.92 0.94 0.93 200
-The two metrics, precision and recall, are best understood by producing a confusion matrix as shown in figure 9 which facilitates assessment of the performance and errors of the CNN model.
+The two metrics, precision and recall, are best understood by producing a confusion matrix, as shown in figure 9 which facilitates assessment of the performance and errors of the CNN model.
 It is evident from the matrix that the model has exhibited strong performance in classifying the scattering events $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O, with high accuracy.
@@
 The scheme of the anode plane segmentation of the SAT-TPC in the simulation model has been varied to 1 mm $\times$ 1 mm and 3 mm $\times$ 3 mm pixels.
-The calculation has been carried out with the same class of events, discussed in the previous section.
+The calculation has been carried out with the same class of events discussed in the previous section.
 The classification performance for the two segmentation schemes in terms of the three metrics is given in table 3.
@@
 This suggests that finer segmentation enhances spatial resolution, allowing the CNN to distinguish subtle differences in track morphology.
-Table 3: Classification report for segmentation schemes 1 mm $\times$ 1 mm and 3 mm $\times$ 3 mm
+Table 3: Classification report for segmentation schemes 3 mm $\times$ 3 mm and 1 mm $\times$ 1 mm
 readout segmentation event classification task Precision Recall F1 Support
-1 mm $\times$ 1 mm direct decay 0.62 0.67 0.64 200
+3 mm $\times$ 3 mm direct decay 0.62 0.67 0.64 200
 sequential decay 0.64 0.57 0.61 200
@@
 $\alpha$+16O scattering 0.80 0.86 0.83 200
-3 mm $\times$ 3 mm direct decay 0.90 0.82 0.86 200
+1 mm $\times$ 1 mm direct decay 0.90 0.82 0.86 200
 sequential decay 0.83 0.90 0.87 200
 $\alpha$+40Ar scattering 0.99 1.00 1.00 200
-$\alpha$+12C 0.99 1.00 0.99 200
+$\alpha$+12C scattering 0.99 1.00 0.99 200
 $\alpha$+16O scattering 0.99 0.98 0.99 200
@@
 4 Summary & Conclusion
-The classification of Hoyle state decay events and the background elastic scattering events forms the central theme of the present study, showcasing the application of supervised machine learning, particularly CNN, for accurate segregation of the relevant class of events.
-The work has highlighted the challenges of distinguishing between direct decay, sequential decay, and the background scattering events $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O in the context of using 30 MeV $\alpha$-beam and a prototype SAT-TPC, filled with active gas target Ar + CO2.
-We have performed Monte Carlo simulations to study the nuclear interactions leading to formation of the Hoyle state, following excitation through inelastic scattering of $\alpha$ from the active target nucleus 12C and the relevant elastic scattering of $\alpha$ from other target nuclei 40Ar, 12C and 16O.
-The decay kinematics have been analyzed in a boosted frame, allowing the extraction of energy and angular distributions of the emitted $\alpha$-particles.
+The classification of Hoyle state decay events and background elastic scattering events forms the central theme of the present study, showcasing the application of a supervised machine learning technique, specifically a CNN, for accurate segregation of the relevant class of events.
+The work has highlighted the challenges of distinguishing between direct decay, sequential decay, and background scattering events $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O in the context of using 30 MeV $\alpha$-beam and a prototype SAT-TPC, filled with active gas target Ar + CO2.
+We have performed Monte Carlo simulations to study the nuclear interactions leading to formation of the Hoyle state, following excitation through inelastic scattering of $\alpha$ from the active target nucleus 12C and the relevant elastic scattering of $\alpha$ from the target nuclei 40Ar, 12C and 16O.
+The decay kinematics have been analyzed in a boosted frame, allowing for the extraction of energy and angular distributions of the emitted $\alpha$-particles.
 The same has been done for the elastic scattering events to study the energy and angular distributions of the scattered products.
-The primary ionization by the beam and the scattering and decay products has been simulated in Geant4 using a low-energy physics list that has been used to reconstruct the event tracks in the SAT-TPC active volume, convoluted with diffusion parameter and projected on two-dimensional readout with segmentation 2 mm $\times$ 2 mm.
+The primary ionization by the beam and the scattering and decay products has been simulated in Geant4 using a low-energy physics list that has been used to reconstruct the event tracks in the SAT-TPC active volume, convoluted with a diffusion parameter, and projected on 2D readout with segmentation of 2 mm $\times$ 2 mm.
 Using the VGG-16 CNN architecture, the model has classified the events on the basis of the track morphology images.
-Regularization technique, such as dropout, has been employed to prevent overfitting and enhance the robustness of the model, achieving high precision, recall, and F1 scores across all categories.
+Dropout was employed as a regularization technique to prevent overfitting and enhance the robustness of the model, achieving high precision, recall, and F1 scores across all categories.
 Despite strong classification performance, the study has identified specific trends in misclassifications, particularly among the decay events.
@@
 These results have proved the efficiency of the model in distinguishing Hoyle state decays from the background events while highlighting areas for further optimization, such as refining feature representations and detector configurations to enhance separation between closely related scattering categories.
-This work has demonstrated the utilization of CNN in analyzing experimental data of low-energy nuclear physics experiments performed with SAT-TPC, providing a framework for advancing robust event classification in nuclear and astrophysical research.
+In the future, the size of the dataset will be increased well above the current choice of 1000 events/class to enhance training diversity and classification performance, especially for the more ambiguous decay modes.
+Additionally, since the CNN is trained on Monte Carlo data, further investigation will be needed to evaluate its robustness and performance on real detector data, including domain adaptation or transfer learning approaches.
+This work has demonstrated the utilization of a CNN in analyzing experimental data of low-energy nuclear physics experiments performed with the SAT-TPC, providing a framework for advancing robust event classification in nuclear and astrophysical research.
 Future work will explore advanced architectures, such as ResNet, and transformer-based models, to further enhance classification performance.
@@
 Tapan Kumar Rana and Dr.
-Santu Manna of the Variable Energy Cyclotron Center, and Prof.
-Satyaki Bhattacharya of Saha Institute of Nuclear Physics Kolkata, India, for valuable discussion and suggestions.
+Santu Manna of the Variable Energy Cyclotron Center, Kolkata and Prof.
+Satyaki Bhattacharya of Saha Institute of Nuclear Physics, Kolkata for valuable discussions and suggestions.